OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x-8

OpenStudy (anonymous):

y=1/2x-8, P(-2,6) solving for dy/dx y=1/2x-8 substitute -2 to x y=1/2(-2)-8 y=(1/-4)-8 y=-33/4 since dy/dx=m m=slope m=-33/4 using point slope form equation of the line y_2-y_1=m(x_2-x_1) let x_2 =x x_1=-2 y_2=y y_1=6 y-6=-33/4(x+2) y-6=(-33/4x)+(-33/2) (-33/4x)+(y)-(6)-(-33/4)=0 (-33/4x)+y-(57/4)=0 <<<<<<<equation of ur tangent line or parallel line